e. The cost of preferred stock 2 marks f. The weighted average cost of capital 7 manks

(20 marks) Question 2 on company has asked its chief financial officer to measure the of each specific form of capital as well as its weighted average cost of asured using the following weights A construction cost capital. The weighted average cost is mea term debt, 10% preferred stock and 50% common stock equity. The company's tax rate is 40%. Debt Company selles $980, a 10 year, $1,000 par value bond that pays a 10% coupon rate annually. Floatation cost is $35 of the per value and the bond is sold at discount of $20 per bond, Preferred Stock: 8% preferred stock with par value of $100 is sold at $65 per share. In addition, the company has to pay an underwriting fee of $2 per share Common stock: The current market value of the common stock is $50 per share. Next year, the dividend (2020) is $4 per share. Its dividend payments which have been approximately 60% of earnings per share in each of the last five years If the company decides to issue a new common equity, then the company must pay the the underpriced costs of $5 per share and floatation costs of $3 per share Based on the information above, Calculate: The before tax cost of debt (5 marks) a. b. The after tax cost of debt (2 marks) c. The cost of common stock (2 marks0 d. The cost of new common stock (2 Marks) LIMKOKWING UNIVERSITY OF CREATIVE TECHNOLOGY
e. The cost of preferred stock 2 marks f. The weighted average cost of capital 7 manks
Answer #1

Before tax cost of debt = C+(F-P)/n /((f+P)/2) = (100+((1000-(980-35))/10))/((1000+(980-35))/2) = 10.85%

After tax cost of debt = before tax cost of debt *(1-t) = 10.85%*(1-40%) = 6.51%

Cost of common stock = D1/P = 4/50 = 8%

Cost of new common stock = D1/(P-U-F) = 4/(50-5-3) = 9.52%

Cost of preferred stock = D/P-U = 8/(65-2) = 12.70%

Weighted average cost of capital = (weight of debt * cost of debt)+(weight of preferred stock * cost of preferred stock)+(weight of common stock * cost of common stock) = (40%*6.51%)+(10%*12.70%)+(50%*9.52%) = 8.63%